# Towards Socially and Morally Aware RL agent: Reward Design With LLM

### v1

@@ 1 Introduction
Reinforcement Learning (RL) is is widely applied in decision-making problems. An agent, the AI system, is trained to find an optimal policy towards satisfying certain objective by maximizing a reward signal by when interacting with the environment through trial and error Sutton and Barto (2018).
The learnt policy may have issues aligning with human valuesHendrycks et al. (2021a) and may cause side effectsAlizadeh Alamdari et al. (2022), and the exploration in finding this policy may be inefficient or unsafeAmodei et al. (2016). As it is difficult to manually specify reward signals for all the things the agent should do and not do while pursuing its goal. Prior approaches to tackle these problems includes using human demonstrationHadfield-Menell et al. (2016), human intervention Saunders et al. (2018) and using language models Hendrycks et al. (2021b); Hafner et al. (2023).
@@ 2 Related Work
Avoiding side effects While pursuing it’s objective, the agent’s interaction of the environment can cause unrelated side effects that are undesired. One approach Alizadeh Alamdari et al. (2022) to avoid negative side effects is by augmenting the reward function to consider other agent’s future value function. Moreover, certain side effect may be not only undesired but dangerous or even catastrophic. This makes the trial and error exploration become unsafe.
Safe Exploration To tackle the previously mentioned issue, Saunders et al. (2018) approach to safe exploration uses human intervention to train a model-free RL agent to learn while avoiding all catastrophic actions in Atari games. At every timestep, a human observes the current state s and the agent’s proposed action a. If the action leads to catastrophic consequence, the human sends a safe action to the environment instead and replaces the new reward with a penalty reward. The difference between safe exploration and avoiding side effect is that in training, the agent decreases the probability and frequency of hitting a negative effect, rather than learning that the effect is negative as the agent explores. This current work investigates whether a language model can be used in place of a human observer in a environment that simulates moral decisions and social norm decisions in real life.
Large Language Model as Reward Kwon et al. (2023) shows the efficiency of leveraging Large Language Model (LLM)’s knowledge as reward signal by prompting the language model to mark each trajectory as good or bad. Wu et al. (2023) further explores the use of LLM by feeding it knowledge of Atari games and prompts it to generate a positive or negative reward that reflects whether an action leads to winning the game. The paper showed improvement in exploration efficiency. Moreover, Hendrycks et al. (2021b) shows that language model can be used to guide RL agents in taking more morally acceptable actions in text-based games.
@@ 3.2 Approach
  • Act as human intervention Saunders et al. (2018) at step level. The objective is to first ensure the exploration is efficient and safe, and second determine the nature of side effects, whether it is positive or negative or neutral. The language model is prompted to generate a reward for all encountered item with a predefined consequence or event. The prompt used is "You are a socially and morally aware person. Please rank the action ${action}$ from -10 to 10, from most undesired to desired. Label your answer as "reasoning:" and "result:" ". The last sentence is replaced with "most appropriate and inappropriate" for experiment 3.3.3. ${action}$ is replaced with the string that describes the consequence or event of the item.
@@ 3.3.3 Public vs Private sphere
The world (figure 8) is designed to test language model’s understanding of social norms. Social normsBicchieri et al. (2023) are context-dependent and ambiguous. In this experiment, there are 3 items on the map: shower, hamburger and cloth with their respective event taking a shower, eating food and changing cloth 8. The appropriateness for each of the items is different depending on whether the event is done is public or not. The goal is to reach for the exit door. The prompt for the language model is: "Consider you as someone in a public space. Please rank the appropriateness of the following action "taking a shower" on a scale of -10 to 10, from least to most appropriate. Label your answer as: "reasoning: " and "result number: number. "". The langugage model used in this experiment is OpenAI’s GPT3.5 model. The human annotated score in appendix A agrees with the language model’s result except for the "eating food in public" event. This is because the language model considers eating in a restaurant where some people surveyed considered eating in public transport and some considered eating in a restaurant. This deviation in the understanding is reflect in the large standard deviation.

### v2

@@ 1 Introduction
Reinforcement Learning (RL) is is widely applied in decision-making problems. An agent, the AI system, is trained to find an optimal policy towards satisfying certain objective by maximizing a reward signal by when interacting with the environment through trial and error [10].
The learnt policy may have issues aligning with human values[6] and may cause side effects[1], and the exploration in finding this policy may be inefficient or unsafe[2]. As it is difficult to manually specify reward signals for all the things the agent should do and not do while pursuing its goal. Prior approaches to tackle these problems includes using human demonstration[4], human intervention [9] and using language models [7, 5].
@@ 2 Related Work
Avoiding side effects While pursuing it’s objective, the agent’s interaction of the environment can cause unrelated side effects that are undesired. One approach [1] to avoid negative side effects is by augmenting the reward function to consider other agent’s future value function. Moreover, certain side effect may be not only undesired but dangerous or even catastrophic. This makes the trial and error exploration become unsafe.
Safe Exploration To tackle the previously mentioned issue, [9] approach to safe exploration uses human intervention to train a model-free RL agent to learn while avoiding all catastrophic actions in Atari games. At every timestep, a human observes the current state s and the agent’s proposed action a. If the action leads to catastrophic consequence, the human sends a safe action to the environment instead and replaces the new reward with a penalty reward. The difference between safe exploration and avoiding side effect is that in training, the agent decreases the probability and frequency of hitting a negative effect, rather than learning that the effect is negative as the agent explores. This current work investigates whether a language model can be used in place of a human observer in a environment that simulates moral decisions and social norm decisions in real life.
Large Language Model as Reward [8] shows the efficiency of leveraging Large Language Model (LLM)’s knowledge as reward signal by prompting the language model to mark each trajectory as good or bad. [11] further explores the use of LLM by feeding it knowledge of Atari games and prompts it to generate a positive or negative reward that reflects whether an action leads to winning the game. The paper showed improvement in exploration efficiency. Moreover, [7] shows that language model can be used to guide RL agents in taking more morally acceptable actions in text-based games.
@@ 3.2 Approach
  • Act as human intervention [9] at step level. The objective is to first ensure the exploration is efficient and safe, and second determine the nature of side effects, whether it is positive or negative or neutral. The language model is prompted to generate a reward for all encountered item with a predefined consequence or event. The prompt used is "You are a socially and morally aware person. Please rank the action ${action}$ from -10 to 10, from most undesired to desired. Label your answer as "reasoning:" and "result:" ". The last sentence is replaced with "most appropriate and inappropriate" for experiment 3.3.3. ${action}$ is replaced with the string that describes the consequence or event of the item.
@@ 3.3.3 Public vs Private sphere
The world (figure 8) is designed to test language model’s understanding of social norms. Social norms[3] are context-dependent and ambiguous. In this experiment, there are 3 items on the map: shower, hamburger and cloth with their respective event taking a shower, eating food and changing cloth 8. The appropriateness for each of the items is different depending on whether the event is done is public or not. The goal is to reach for the exit door. The prompt for the language model is: "Consider you as someone in a public space. Please rank the appropriateness of the following action "taking a shower" on a scale of -10 to 10, from least to most appropriate. Label your answer as: "reasoning: " and "result number: number. "". The langugage model used in this experiment is OpenAI’s GPT3.5 model. The human annotated score in appendix A agrees with the language model’s result except for the "eating food in public" event. This is because the language model considers eating in a restaurant where some people surveyed considered eating in public transport and some considered eating in a restaurant. This deviation in the understanding is reflect in the large standard deviation.